OpenStudy (anonymous):

Find center-radius: 8x^2+8x+8y^2-64y-382=0

OpenStudy (danjs):

recall - a circle with a center (h,k) and radius r , can look like (x - h)^2 + (y - k)^2 = r^2 for things that look like what they gave you , usually for a book problem it is completing the square with nice whole numbers prolly

OpenStudy (danjs):

i would divide by 8 first \[x^2 + x + y^2 - 8y = 382/8\] then just turn those into a perfect square , for ax^2 + bx + ? , you can take half of b, and square that

OpenStudy (danjs):

x^2 + x + (1/2)^2 + y^2 - 8y + (-4)^2 = 382/8 + (1/2)^2 + (-4)^2

OpenStudy (danjs):

(x + 1/2)^2 + (y - 4)^2 = 382/8 + 1/4 + 16 there is your (h,k) , simplify the other side to see r^2 to get the radius